Part B 2K(s) + Br2(l) → 2 KBr(s) Express your answer using four significant figures.
media%2F761%2F761ddad9-aed5-4692-a7a0-6f
Answer #1

15.83 g of Br2 (molar mass 159.808 g/mol) corresponds to

\frac{\text { 15.83 g }}{\text { 159.8 g/mol}}=\text { 0.09906 mol}

1 mole of Br2 gives 2 moles of KBr.

0.09906 moles of Br2 gives 2 \times \text { 0.09906 mol}=\text { 0.198123 mol} of KBr.

The molar mass of KBr is 119 g/mol.

Mass of KBr formed =\frac{\text { 0.198123 mol}}{ \text { 119 g/mol }} = \text { 0.001665 g}